OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x to the 6th power over x to the second power. Explain in your own words how you can simplify x to the 6th power over x to the second power without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

jhonyy9 (jhonyy9):

x^6 ---- = x^(6-2) =x^4 x^2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

So my answer would be x^4?